NOAA SATOPS Morning Report: Wednesday, October 31, 2012


SATOPS Morning Report: Wednesday, October 31, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 305

GOES
GOES-15 CDA support moved to Fairbanks from 21:12 on DOY302 to 21:03z on DOY304.
GOES-12 DOY304 intermittent frame breaks due to high winds over WCDA

All other GOES operations were nominal over the past 120 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
POES Operations suspended at WCDA(due to hurricane Sandy) from 14:00z on DOY302 through 22:38z on DOY304. N19 call up supports were taken at IJPS Svalbard to avoid data loss. N19 LACs over IJPS Svalbard not supported and therefore lost. Call up supports taken at FCDA for N15, N16, N17, and N18 in order to minimize data loss.

N17 DOY302 23:24z G2B, G1A and G1B not recovered.
N18 DOY303 07:52z G5A G4B and G1A not recovered.
N15 DOY303 09:22z G4A G1A and G4B not recovered
N19 DOY303 05:09z G5B from SVL not recovered due to a N19 stx3 problem. Under investigation.
N19 DOY304 10:22z FCDA L5A recorded was stopped and repositioned in order to chase G1B to be sent to Svalbard at scheduled on STX3 which has a problem. L5A only received 2059 of 4140 frames

All other POES operations were nominal over the past 120 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 120 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
Jason-2 Operations suspended at WCDA due to hurricane Sandy from 14:00z on DOY302 through 22:38z on DOY304. Jason-2 call up support were taken at FCDA in order to minimize data latency. No data lost.

All other JASON-2 operations were nominal over the past 120 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
DOY301 ESPC did not receive the 0240Z WCONUS image on the primary Gini L1 West. Under investigation.

DOY303 The primary West GINI hung on the GOES-west 0245Z, 0300Z, 0330Z images because of frame breaks and a GOES ground system problem

DOY305 AFWA (Air Force Weather Agency) notified ESPC that the 05z North and South Cloud data will be missed due to system issues at AFWA.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.